# Extremal random matrices with independent entries and matrix superconcentration inequalities

### v1

@@ 3. The symmetric case
whenever $b_{ii}=\sigma_{*}$ and $\xi_{ii}\sim N(0,1)$ for all $i$ (as then $X_{ii}=\sqrt{2}\sigma_{*}\xi_{ii}$). Moreover, it follows from [17, Corollary 3.2] that the tail bounds of both Theorems 2.4 and 3.5 cannot be improved in general for $t\to\infty$ (even in models where $b_{ii}=0$ for all $i$).
@@ 3.2. Small deviations
With the extremum principle in hand, we can now repeat the arguments of section 2.2 in the symmetric case. The main difference is that now $Y$ is a GOE matrix rather than a GUE matrix. However, a suitable moment estimate for GOE matrices was also obtained by Ledoux [16, Theorem 8].
@@ 3.3. Large deviations
We clearly have $\mathop{\mathrm{tr}}e^{tY}\leq e^{t\lambda_{1}(Y)}$, where $\lambda_{1}(Y)$ denotes the largest eigenvalue of $Y$. It is shown in [8, Theorem 2.11] that $\mathbf{E}[\lambda_{1}(Y)]\leq 2\sqrt{d}$. On the other hand, the Gaussian log-Sobolev inequality implies by [13, eq. (5.8)] and [8, §2.2] that $\lambda_{1}(Y)$ is a $\sqrt{2}$-subgaussian random variable, that is, that
@@ 4.3. Large deviations
We clearly have $\mathop{\mathrm{tr}}e^{t(YY^{*})^{1/2}}\leq e^{t\|Y\|}$. It is shown in [8, Theorem 2.13] that $\mathbf{E}\|Y\|\leq\sqrt{d_{1}}+\sqrt{d_{2}}$. The Gaussian log-Sobolev inequality implies by [13, eq. (5.8)] and [8, §2.2] that $\|Y\|$ is a $1$-subgaussian random variable, that is, that
@@
This yields the following.
Proposition 4.13.
For $X$ as in Model 1.3 with $n\leq m$, we have
for all $\varepsilon\geq 0$.
Proof.
Suppose first that $\sigma_{*}=1$, and let $d_{1}=\lceil\sigma_{1}^{2}\rceil$ and $d_{2}=\lceil\sigma_{2}^{2}\rceil$. Then
@@
$\leq\mathbf{E}[\mathop{\mathrm{tr}}\cosh(t(YY^{*})^{1/2})]\leq\mathbf{E}\big{[}\mathop{\mathrm{tr}}e^{t(YY^{*})^{1/2}}\big{]},$
@@
for all $\varepsilon\geq 0$. This concludes the proof for $\sigma_{*}=1$. For general $\sigma_{*}$, it suffices to apply the above bound to the random matrix $\frac{X}{\sigma_{*}}$. ∎
Theorem 1.6 follows readily.
Proof of Theorem 1.6.
The conclusion follows immediately from Proposition 4.13 using that $\sqrt{\sigma_{1}^{2}+\sigma_{*}^{2}}+\sqrt{\sigma_{2}^{2}+\sigma_{*}^{2}}\leq\sigma_{1}+\sigma_{2}+\frac{\sigma_{*}^{2}}{2\sigma_{1}}+\frac{\sigma_{*}^{2}}{2\sigma_{2}}\leq\sigma_{1}+\sigma_{2}+\sigma_{*}$. ∎
@@ 5. Moment estimates for Wishart matrices
A moment estimate for complex Wishart matrices was previously obtained by Ledoux in [14, p. 201]. However, the constants in Ledoux’ estimate depend in an unspecified manner on the aspect ratio $c$, making it unsuitable for the purposes of this paper. [^3] The argument of [14, p. 201] also contains a further issue, that the recursion in eq. (29) of that paper does not imply the inequality for $b_{p}$ claimed subsequently. The difficulty in the proofs of Theorems 5.1 and 5.2 is to obtain bounds that have optimal dependence on $c$, which requires a more delicate understanding of the structure of the associated moment recursions.
@@
On the other hand, if we let $n\to\infty$ with $p,c$ fixed, it is classical [18, p. 368] that both $n^{-p}\mathbf{E}[\mathop{\mathrm{tr}}(YY^{*})^{p}]$ and $n^{-p}\mathbf{E}[\mathop{\mathrm{tr}}(ZZ^{*})^{p}]$ converge to the $p$-moment $\chi_{p}^{c}$ of the Marchenko–Pastur distribution. By using the explicit formula for its generating function [18, p. 205], the Darboux method [23, Theorem 5.11] yields
@@
The remainder of this section is organized as follows. In section 5.1, we recall the recursive formulas of [10, 7] for the moments of complex and real Wishart matrices. Let us emphasize at the outset that the recursive formula for real Wishart moments involves complex Wishart moments, so that we must consider the latter even if one is ultimately only interested in real Wishart matrices. We then prove Theorem 5.1 in section 5.2, and finally prove Theorem 5.2 in section 5.3.
@@ 5.1. Moment recursions
The proofs of the main results of this section are based on recursive formulas for $A_{p},A_{p}^{\prime},B_{p}$ that we presently recall. The following recursive formula for $A_{p}$ was obtained by Haagerup and Thorbjørnsen [10, Theorem 8.2].
@@
The reason that we consider the modified moments $A_{p}^{\prime}$ is that they appear in the following moment recursion for $B_{p}$ due to Cunden et al. [7, Theorem 3.5] (the values of $B_{1},B_{2}$ stated here are readily obtained by a straightforward computation).
@@ 5.3. Real Wishart moments
Taking inspiration from [16], we define
@@ Acknowledgment
This work was supported in part by NSF grants DMS-1856221 and DMS-2054565.

### v2

@@ 3. The symmetric case
whenever $b_{ii}=\sigma_{*}$ and $\xi_{ii}\sim N(0,1)$ for all $i$ (as then $X_{ii}=\sqrt{2}\sigma_{*}\xi_{ii}$). Moreover, it follows from [LT91, Corollary 3.2] that the tail bounds of both Theorems 2.4 and 3.5 cannot be improved in general for $t\to\infty$ (even in models where $b_{ii}=0$ for all $i$).
@@ 3.2. Small deviations
With the extremum principle in hand, we can now repeat the arguments of section 2.2 in the symmetric case. The main difference is that now $Y$ is a GOE matrix rather than a GUE matrix. However, a suitable moment estimate for GOE matrices was also obtained by Ledoux [Led09, Theorem 8].
@@ 3.3. Large deviations
We clearly have $\mathop{\mathrm{tr}}e^{tY}\leq e^{t\lambda_{1}(Y)}$, where $\lambda_{1}(Y)$ denotes the largest eigenvalue of $Y$. It is shown in [DS01, Theorem 2.11] that $\mathbf{E}[\lambda_{1}(Y)]\leq 2\sqrt{d}$. On the other hand, the Gaussian log-Sobolev inequality implies by [Led01, eq. (5.8)] and [DS01, §2.2] that $\lambda_{1}(Y)$ is a $\sqrt{2}$-subgaussian random variable, that is, that
@@ 4.3. Large deviations
We clearly have $\mathop{\mathrm{tr}}e^{t(YY^{*})^{1/2}}\leq e^{t\|Y\|}$. It is shown in [DS01, Theorem 2.13] that $\mathbf{E}\|Y\|\leq\sqrt{d_{1}}+\sqrt{d_{2}}$. The Gaussian log-Sobolev inequality implies by [Led01, eq. (5.8)] and [DS01, §2.2] that $\|Y\|$ is a $1$-subgaussian random variable, that is, that
@@
We can now prove Theorem 1.6.
Proof of Theorem 1.6.
Suppose that $\sigma_{*}=1$, and let $d_{1}=\lceil\sigma_{1}^{2}\rceil$, $d_{2}=\lceil\sigma_{2}^{2}\rceil$. Then
@@
$\leq\mathbf{E}[\mathop{\mathrm{tr}}\cosh(t(YY^{*})^{1/2})]\leq\mathbf{E}\big{[}\mathop{\mathrm{tr}}e^{t(YY^{*})^{1/2}}\big{]}$
@@
for all $\varepsilon\geq 0$.
For general $\sigma_{*}$, applying the above bound to the random matrix $\frac{X}{\sigma_{*}}$ yields
for all $\varepsilon\geq 0$. To conclude the proof, it remains to use that $\sqrt{\sigma_{1}^{2}+\sigma_{*}^{2}}+\sqrt{\sigma_{2}^{2}+\sigma_{*}^{2}}\leq\sigma_{1}+\sigma_{2}+\frac{\sigma_{*}^{2}}{2\sigma_{1}}+\frac{\sigma_{*}^{2}}{2\sigma_{2}}\leq\sigma_{1}+\sigma_{2}+\sigma_{*}$. ∎
@@ 5. Moment estimates for Wishart matrices
A moment estimate for complex Wishart matrices was previously obtained by Ledoux in [Led04, p. 201]. However, the constants in Ledoux’ estimate depend in an unspecified manner on the aspect ratio $c$, making it unsuitable for the purposes of this paper. [^4] The argument of [Led04, p. 201] also contains a further issue, that the recursion in eq. (29) of that paper does not imply the inequality for $b_{p}$ claimed subsequently. The difficulty in the proofs of Theorems 5.1 and 5.2 is to obtain bounds that have optimal dependence on $c$, which requires a more delicate understanding of the structure of the associated moment recursions.
@@
On the other hand, if we let $n\to\infty$ with $p,c$ fixed, it is classical [NS06, p. 368] that both $n^{-p}\mathbf{E}[\mathop{\mathrm{tr}}(YY^{*})^{p}]$ and $n^{-p}\mathbf{E}[\mathop{\mathrm{tr}}(ZZ^{*})^{p}]$ converge to the $p$-moment $\chi_{p}^{c}$ of the Marchenko–Pastur distribution. By using the explicit formula for its generating function [NS06, p. 205], the Darboux method [Wil06, Theorem 5.11] yields
@@
The remainder of this section is organized as follows. In section 5.1, we recall the recursive formulas of [HT03, CMSV16] for the moments of complex and real Wishart matrices. Let us emphasize at the outset that the recursive formula for real Wishart moments involves complex Wishart moments, so that we must consider the latter even if one is ultimately only interested in real Wishart matrices. We then prove Theorem 5.1 in section 5.2, and finally prove Theorem 5.2 in section 5.3.
@@ 5.1. Moment recursions
The proofs of the main results of this section are based on recursive formulas for $A_{p},A_{p}^{\prime},B_{p}$ that we presently recall. The following recursive formula for $A_{p}$ was obtained by Haagerup and Thorbjørnsen [HT03, Theorem 8.2].
@@
The reason that we consider the modified moments $A_{p}^{\prime}$ is that they appear in the following moment recursion for $B_{p}$ due to Cunden et al. [CMSV16, Theorem 3.5] (the values of $B_{1},B_{2}$ stated here are readily obtained by a straightforward computation).
@@ 5.3. Real Wishart moments
Taking inspiration from [Led09], we define
@@ Acknowledgment
This work was supported in part by NSF grants DMS-1856221, DMS-2054565, and DMS-2347954. We thank the referee for helpful suggestions.